OpenStudy (anonymous):

\[f(x) =\begin{cases}0 & x = \text{rational}\\1 & x = \text{irrational}\end{cases}\] Find \[\int_0^1f(x)dx\]

terenzreignz (terenzreignz):

You want to integrate a non-integrable function? D:

OpenStudy (anonymous):

Hmmmm... Why not?

terenzreignz (terenzreignz):

Why it's not integrable?

OpenStudy (anonymous):

Yes, why is it not integrable?

terenzreignz (terenzreignz):

Upper and lower sums are never equal.

OpenStudy (anonymous):

Are you sure that it is NOT integrable?

terenzreignz (terenzreignz):

Quite... it is frequently used as an example of something which isn't integrable. Let me show you what I mean... when you divide it into sub-intervals, each sub-interval will inevitably contain both rational and irrational numbers, right?

OpenStudy (anonymous):

But can't we select a VERY small interval which contains only irrational numbers?

terenzreignz (terenzreignz):

No, that's the beauty of it, there is always a rational number between any two irrational numbers, and vice versa, that's the density property :D

OpenStudy (anonymous):

Then what about this? "Let c be an irrational number in [0,1] and ϵ > 0. Then there is a natural number n such that 1/n < ϵ. If we choose δ small enough that the interval (c-δ, c+δ) contain no rational numbers with denominator less than n, then it follows that for x in this interval, we have f(x)-f(c)| = |f(x)| ≤ 1/n < ϵ" Extracted from http://www.math.harvard.edu/~ctm/home/text/class/harvard/112/02/html/home/solns/sol11.pdf

terenzreignz (terenzreignz):

While I don't really see what that proves, but suppose we have two irrational numbers, \(\alpha\) and \(\beta\) (or real numbers, it don't matter ;) ) Then, suppose \(\beta > \alpha\) (one of them has to be bigger than the other of course) Then, \[\beta - \alpha > 0\] There must exist a positive integer that is bigger than \(\LARGE \frac1{\beta-\alpha}\) , let that integer be n. \[\Large n > \frac{1}{\beta - \alpha}\]\[\Large n\beta - n\alpha > 1\]This means there is an integer m between \(n\alpha \) and \(n\beta\) since their difference is more than 1. \[\Large n\alpha < m < n\beta\]\[\Large \alpha < \frac{m}n<\beta\] and there's your rational number.

OpenStudy (anonymous):

"This means there is an integer m between nα and nβ since their difference is more than 1." Why must m be an integer?

terenzreignz (terenzreignz):

If two numbers are more than 1 unit apart, there must be an integer between them.

terenzreignz (terenzreignz):

While I think of how to prove this, just you try to find a counterexample... now let's see...

terenzreignz (terenzreignz):

okay, a rough sketch... first get rid of the trivial case, when one of them is an integer, obviously, if the smaller number \(\alpha\) is an integer, then \(\alpha + 1\) is the integer between \(\alpha\) and \(\beta\). A similar argument, involving \(\beta - 1\) goes if \(\beta\) is an integer. Now, every non-integer is between two integers...

terenzreignz (terenzreignz):

So that \(\Large \alpha < \lceil \alpha \rceil\)

terenzreignz (terenzreignz):

But \[\Large 0<\lceil \alpha \rceil - \alpha <1\] since \(\alpha \) is not an integer.

terenzreignz (terenzreignz):

\[\Large \beta - \alpha > 1 > \lceil \alpha \rceil-\alpha \]\[\Large \implies \beta - \alpha > \lceil\alpha\rceil-\alpha\implies \beta> \lceil\alpha\rceil\] \[\Large \implies \alpha < \lceil\alpha\rceil< \beta\]

terenzreignz (terenzreignz):

Thus there must be an integer between two numbers if their difference is more than 1.

OpenStudy (anonymous):

I got the part that "when you divide it into sub-intervals, each sub-interval will inevitably contain both rational and irrational numbers". Please move on.

terenzreignz (terenzreignz):

Okay, well, once you divide the interval into sub-intervals, the upper sum would always be 1, right? Since in each interval, the highest possible value f could take is 1, since each interval has irrationals...

OpenStudy (anonymous):

Yes.

terenzreignz (terenzreignz):

On the other hand, the lower sum would always be zero since the lowest possible value f could take is 0, since each interval would have rationals.

OpenStudy (anonymous):

Wait... lower sum or lower limit?

terenzreignz (terenzreignz):

lower sum, lower limit may or may not refer to the same thing, I'm not aware :)

OpenStudy (anonymous):

Ok...

terenzreignz (terenzreignz):

Well, the upper sum is always 1, the lower sum is always 0, the upper sum and lower sum must always approach each other (so to speak) for the function to be Riemann Integrable...

OpenStudy (anonymous):

Why in the pdf does it mention that it is integrable?

terenzreignz (terenzreignz):

Interesting. Lebesgue integral by any chance?

OpenStudy (anonymous):

Would you please elaborate it?

terenzreignz (terenzreignz):

I don't know it myself, I just saw it in passing that the Dirichlet function (a function similar to this one) is *not* Riemann Integrable, but it is Lebesgue Integrable. That's all I know on the matter :>

OpenStudy (anonymous):

Alright. Thanks for your help!

OpenStudy (anonymous):

Do you know if anyone might know this?

terenzreignz (terenzreignz):

Try the big guys :D

OpenStudy (anonymous):

Feel free to tag them :)

terenzreignz (terenzreignz):

You tag them :P

terenzreignz (terenzreignz):

I'd rather not bother them myself XD

OpenStudy (anonymous):

Right, You need to do something if you want to learn, so... @amistre64 @satellite73 @zarkon @hartnn @KingGeorge The questions are 1) Why is \(f(x) =\begin{cases}0 & x = \text{rational}\\1 & x = \text{irrational}\end{cases}\) integrable over [0,1]? 2) How to find \(\int_0^1f(x)dx\)? Any help would be highly appreciated!

terenzreignz (terenzreignz):

<fingers crossing>

OpenStudy (kinggeorge):

First off, \(f(x) =\begin{cases}0 & x = \text{rational}\\1 & x = \text{irrational}\end{cases}\) is not Riemann integrable for the reasons that @terenzreignz gave above. However, it IS Lebesgue integrable. This can be seen by noticing that the set \(\mathbb{Q}\cap[0,1]\) is a countable set, and thus has measure 0. So the set \([0,1]\setminus(\mathbb{Q}\cap[0,1])\), which is just the irrational numbers in the interval [0,1], has measure 1. So by the definition of Lebesgue integral on indicator functions, we have that\[\int_0^1f(x) dx=1.\]